Circular No. 8130

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ee Further to IAUC 8117, W. M. Wood-Vasey, G. Aldering, and P. Nu-

gent report the discovery of an apparent supernova (mag 16.5) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on Mar. 28.20–28.24 UT. The new object is located at α = 11^{h}33^{m}24^{s}.50, δ = −9^{o}59′26′′.4 (equinox 2000.0), which is 1′′ east and 1′′ south of the center of the apparent host galaxy. Additional NEAT magnitudes from observations at Palomar and Haleakala: 2002 Apr. 5.20, [21.6; 2003 Feb. 9.51, [21.0; Mar. 3.58, 17.8; 27.32, 16.5; Apr. 7.34, 16.5; 10.13, 16.5; 11.22, 16.6.

M. Phillips and M. Hamuy, Carnegie Observatories, report that a spectrum (range 360–900 nm) of SN 2003ee, obtained on May 5.10 UT with the Las Campanas Observatory Clay 6.5-m telescope (+ LDSS-2), shows it to be a type-IIn supernova characterized by a blue continuum and strong Balmer lines in emission, at a redshift of 0.021.

K. Krisciunas, National Optical Astronomy Observatory, reports that photometry on Apr. 21.223 UT with the Las Campanas 1-m telescope yields the following magnitudes for SN 2003ee: B = 16.628 ± 0.024, V = 16.548 ± 0.017, I = 16.383 ± 0.024.

SUPERNOVAE 2003dq AND 2003dt Phillips and Hamuy add that a spectrum of SN 2003dq (cf. IAUC

8117), taken on May 5.06 UT, shows it to be a type-II supernova with Balmer lines exhibiting P-Cyg profiles, although the Hα absorption is very weak. The expansion velocity of −7800 km/s, derived from the minimum of Hα (based on a redshift of 0.046, derived for the host galaxy from superimposed narrow emission Balmer lines), suggests that SN 2003dq is several weeks old. A spectrum of SN 2003dt (cf. IAUC 8120), taken on May 5.35, shows it to be a type-Ia supernova, ∼ 3 months past maximum light.

2003ec K. Itagaki, Teppo-cho, Japan, reports (via S. Nakano) that his unfil-

tered CCD images taken on 2001 May 13 and 2002 May 3 with a 0.60-m reflector clearly show the object reported by Chen et al. on IAUC 8128. Phillips and Hamuy also note that a spectrum of 2003ec, taken as above on May 5.33 UT, shows a blue, featureless continuum.

2003 May 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green